OpenStudy (anonymous):

pwease im begging for some help :3 will do what ever u want :P \(e^{i\theta}=\cos\theta+i\sin\theta\) prove that \((e^{i\theta})^n=e^{i n \theta}\) for all n in Z

ganeshie8 (ganeshie8):

Given : \(\large e^{i\theta}=\cos\theta+i\sin\theta \) you want to prove : \((e^{i\theta})^n=e^{in\theta}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its correct if i use exponential properties , but i should use the given equation

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles What you want to show: \[\large e^{in\theta}=\cos n\theta+i\sin n\theta=(\cos \theta+i\sin\theta)^n=\left(e^{i\theta}\right)^n\] Try it by induction, with the base case \(n=1\): \[\large e^{i\theta}=\cos\theta+i\sin\theta=\left(e^{i\theta}\right)^1\] This is true by the definition of \(e^{i\theta}\). Assume equality holds for \(n=k\): \[\large e^{ik\theta}=\cos k\theta+i\sin k\theta=(\cos\theta+i\sin\theta)^k=\left(e^{i\theta}\right)^k\] Show it holds for \(n=k+1\): \[\large e^{i(k+1)\theta}=e^{ik\theta}e^{i\theta}=\cdots=\left(e^{i\theta}\right)^{k+1}\] Fill in the blanks. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

good!it works for n>0 what about n<0 ?

OpenStudy (anonymous):

oh well ty alote @SithsAndGiggles for ur help ! i got it now for \(n=0 \) and \(n<0\)